# Investigating and Enhancing Vision-Audio Capability in Omnimodal Large Language Models

### v1

@@ Abstract
Omnimodal Large Language Models (OLLMs) have shown significant progress in integrating vision and text, but still struggle with integrating vision and audio, often exhibiting suboptimal performance when processing audio queries compared to text queries. This disparity is primarily due to insufficient alignment between vision and audio modalities during training, leading to inadequate attention to visual information when using audio queries. To mitigate this issue, we propose a Self-Knowledge Distillation (Self-KD) training method where the vision-text component of the OLLM serves as the teacher and the vision-audio component as the student. This enables the model to process audio in a manner analogous to its text processing. Our experimental results demonstrate that Self-KD is an effective method for enhancing the vision-audio capabilities of OLLMs by learning from the vision-text components, which subsequently improves the interaction between audio and images and results in improved performance on multimodal tasks.
@@
Rui Hu1thanks: This work was done during an internship at Unisound., Delai Qiu2, Shuyu Wei1, Jiaming Zhang1, Yining Wang2, Shengpeng Liu2, Jitao Sang1thanks: Corresponding author. 1Beijing Jiaotong University 2Unisound AI Technology Co., Ltd. rui.hu@bjtu.edu.cn, jtsang@bjtu.edu.cn
@@ 1 Introduction
Recent years have witnessed significant advancements in large language models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Yang et al. (2024), which have catalyzed the development of multimodal large language models (MLLMs) Wang et al. (2024b); Chen et al. (2024a); Liu et al. (2024); Fang et al. (2024); Chu et al. (2024). This progress marks a paradigm shift in how machines understand and interact with the world, with omnimodal large language models (OLLMs) OpenAI (2024); Fu et al. (2024); Xie and Wu (2024); Fu et al. (2025); Li et al. (2024); InfinigenceAI (2024) emerging as a new frontier. These models, exemplified by GPT-4o, demonstrate advanced capabilities in visual, linguistic, and auditory functionalities, promising more natural and comprehensive interactions.
@@
However, despite these advancements, a critical gap remains in the performance of OLLMs when processing vision-text versus vision-audio inputs. Specifically, OLLMs often exhibit suboptimal performance on vision-audio tasks compared to their vision-language counterparts. For instance, replacing a text question with its audio equivalent can result in contradictory responses from the models. As illustrated in Figure 1, when the text question ‘‘What’s the name of the book on the top of the pile?’’ is posed to Megrez InfinigenceAI (2024), the model accurately responds with ‘‘Ariel’’. However, when the same question is converted into audio, it erroneously answers ‘‘Plays pleasant’’. This inconsistency is prevalent across various OLLMs, indicating that the models exhibit different behaviors when processing vision-text and vision-audio inputs.
@@
(1) We identify and analyze the significant gap in performance between vision-language and vision-audio capabilities in OLLMs, attributed to insufficient alignment and between images and audio.
(2) We propose a Self-KD training framework that leverages the vision-text component to guide the training of the vision-audio component, promoting better alignment and integration of visual and audio information.
(3) We conduct extensive experiments on various models and datasets, demonstrating that Self-KD significantly enhances vision-audio performance compared to conventional vision-audio SFT.
@@ 2.1 Setup
OLLMs. We select three open-source OLLMs, VITA Fu et al. (2024), VITA-1.5 Fu et al. (2025), and Megrez InfinigenceAI (2024) for testing, with parameter sizes of 8×7B, 7B, and 3B, respectively.
@@ 3.1 Attention Weight Analysis
Query tokens pay less attention to vision tokens under audio queries than under text queries: In Figure 4(a), we present how the attention weights from query tokens to image tokens and to themselves vary across different layers in Megrez InfinigenceAI (2024). This reflects how the model processes the input information to prepare for output. Consistent with the findings of Bi et al. (2024); Zhang et al. (2025), we observe that the model’s attention to vision tokens is high in the early layers, regardless of whether the query is text or audio. However, in the middle and later layers of the model, when using an audio query, the attention weights from the query tokens to the vision tokens are consistently lower than those with a text query. In contrast, the model focuses more on the query token itself. This suggests that the model may struggle to effectively integrate audio and visual information in the later layers, leading to the inferior performance on vision-audio tasks compared to vision-text tasks.
@@ 5.1 Experimental Setup
Models. We select the InternVL2 series Chen et al. (2024b) and Qwen2VL series Wang et al. (2024a) as our base models due to their excellent performance and the availability of multiple sizes. Following Li et al. (2024); Chu et al. (2024), we use the Whisper-large-v3 model Radford et al. (2023) as the audio encoder and a one-layer MLP as the projector to convert audio features to LLM embeddings.
@@ 6 Related Works
Omnimodal Large Language Models. Recent advancements in multimodal large models have primarily focused on Vision-Language Models, e.g., CLIP Radford et al. (2021), followed by models such as Intern-VL Chen et al. (2024a), which use MLPs to integrate vision encoders and LLMs for enhanced semantic alignment. Audio-Language Models, like Qwen-Audio Chu et al. (2024), combine audio encoders with LLMs to directly map audio signals to text. Recently, Omnimodal Large Language Models (OLLMs) have emerged, integrating vision, audio, and text by aligning their encoders during training for end-to-end processing. Models such as VITA Fu et al. (2024, 2025), Mini-Omni2 Xie and Wu (2024), MiniCPM-o MiniCPM-o Team (2025), and Baichuan-Omni Li et al. (2024) have demonstrated strong multimodal performance.
@@
Knowledge Distillation in MLLMs. Knowledge distillation Hinton (2015) has recently been applied to multimodal large language models (MLLMs). For example, LLaVA-MoD Shu et al. (2024) and LLaVA-KD Cai et al. (2024) use knowledge distillation to transfer the performance of large teacher models to smaller student models. This paper proposes a self-knowledge distillation method, dividing the same model into teacher and student components to bring the vision-audio capabilities of OLLMs closer to their vision-text capabilities.

### v2

@@ Abstract
Omnimodal Large Language Models (OLLMs) have shown significant progress in integrating vision and text, but still struggle with integrating vision and audio, often exhibiting suboptimal performance when processing audio queries compared to text queries. This disparity is primarily due to insufficient alignment between vision and audio modalities during training, leading to inadequate attention to visual information when using audio queries. To mitigate this issue, we propose a Self-Knowledge Distillation (Self-KD) training method where the vision-text component of the OLLM serves as the teacher and the vision-audio component as the student. This enables the model to process audio in a manner analogous to its text processing. Our experimental results demonstrate that Self-KD is an effective method for enhancing the vision-audio capabilities of OLLMs by learning from the vision-text components, which subsequently improves the interaction between audio and images and results in improved performance on multimodal tasks [^1] https://github.com/isruihu/Self-KD.
@@
Rui Hu1thanks: Work done during an internship at Unisound., Delai Qiu2, Shuyu Wei1, Jiaming Zhang1, Yining Wang2, Shengpeng Liu2, Jitao Sang1,3thanks: Corresponding author. 1Beijing Key Lab of Traffic Data Analysis and Mining, Beijing Jiaotong University 2Unisound AI Technology Co., Ltd. 3Peng Cheng Lab rui.hu@bjtu.edu.cn
@@ 1 Introduction
Recent years have witnessed significant advancements in large language models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Yang et al. (2024), which have catalyzed the development of multimodal large language models (MLLMs) Wang et al. (2024b); Chen et al. (2024b); Liu et al. (2024); Fang et al. (2024); Chu et al. (2024); Tang et al. (2024). This progress marks a paradigm shift in how machines understand and interact with the world, with omnimodal large language models (OLLMs) Hurst et al. (2024); Team et al. (2023); Fu et al. (2024); Xie and Wu (2024); Fu et al. (2025); Li et al. (2024, 2025a); Chen et al. (2024a); Luo et al. (2025) emerging as a new frontier. These models, exemplified by GPT-4o Hurst et al. (2024), demonstrate advanced capabilities in visual, linguistic, and auditory functionalities, promising more natural and comprehensive interactions.
@@
However, despite these advancements, a critical gap remains in the performance of OLLMs when processing vision-text versus vision-audio inputs. Specifically, OLLMs often exhibit suboptimal performance on vision-audio tasks compared to their vision-language counterparts. For instance, replacing a text question with its audio equivalent can result in contradictory responses from the models. As illustrated in Figure 1, when the text question ‘‘What’s the name of the book on the top of the pile?’’ is posed to Megrez Li et al. (2025a), the model accurately responds with ‘‘Ariel’’. However, when the same question is converted into audio, it erroneously answers ‘‘Plays pleasant’’. This inconsistency is prevalent across various OLLMs, indicating that the models exhibit different behaviors when processing vision-text and vision-audio inputs.
@@
  • We identify and analyze the significant gap in performance between vision-language and vision-audio capabilities in OLLMs, attributed to insufficient alignment and between images and audio.
  • We propose a Self-KD training framework that leverages the vision-text component to guide the training of the vision-audio component, promoting better alignment and integration of visual and audio information.
  • We conduct extensive experiments on various models and datasets, demonstrating that Self-KD significantly enhances vision-audio performance compared to conventional vision-audio SFT.
@@ 2.1 Setup
OLLMs. We select three open-source OLLMs, VITA Fu et al. (2024), VITA-1.5 Fu et al. (2025), and Megrez Li et al. (2025a) for testing, with parameter sizes of 8×7B, 7B, and 3B, respectively.
@@ 3.1 Attention Weight Analysis
Query tokens pay less attention to vision tokens under audio queries than under text queries: In Figure 4(a), we present how the attention weights from query tokens to image tokens and to themselves vary across different layers in Megrez Li et al. (2025a). This reflects how the model processes the input information to prepare for output. Consistent with the findings of Bi et al. (2024); Zhang et al. (2025), we observe that the model’s attention to vision tokens is high in the early layers, regardless of whether the query is text or audio. However, in the middle and later layers of the model, when using an audio query, the attention weights from the query tokens to the vision tokens are consistently lower than those with a text query. In contrast, the model focuses more on the query token itself. This suggests that the model may struggle to effectively integrate audio and visual information in the later layers, leading to the inferior performance on vision-audio tasks compared to vision-text tasks.
@@ 5.1 Experimental Setup
Models. We select the InternVL2 series Chen et al. (2024c) and Qwen2VL series Wang et al. (2024a) as our base models due to their excellent performance and the availability of multiple sizes. Following Li et al. (2024); Chu et al. (2024), we use the Whisper-large-v3 model Radford et al. (2023) as the audio encoder and a one-layer MLP as the projector to convert audio features to LLM embeddings.
@@ 6 Related Works
Omnimodal Large Language Models. Recent advancements in multimodal large models have primarily focused on Vision-Language Models, e.g., CLIP Radford et al. (2021), followed by models such as Intern-VL Chen et al. (2024b), which use MLPs to integrate vision encoders and LLMs for enhanced semantic alignment. Audio-Language Models, like Qwen-Audio Chu et al. (2024), combine audio encoders with LLMs to directly map audio signals to text. Recently, Omnimodal Large Language Models (OLLMs) have emerged, integrating vision, audio, and text by aligning their encoders during training for end-to-end processing. Models such as VITA Fu et al. (2024, 2025), Mini-Omni2 Xie and Wu (2024), MiniCPM-o MiniCPM-o Team (2025), Baichuan-Omni Li et al. (2024, 2025b) and Qwen2.5-Omni Xu et al. (2025) have demonstrated strong multimodal performance.
@@
Knowledge Distillation in MLLMs. Knowledge distillation Hinton (2015); Gu et al. (2024); Wang and Yoon (2021) has recently been applied to multimodal large language models (MLLMs). For example, LLaVA-MoD Shu et al. (2024) and LLaVA-KD Cai et al. (2024) use knowledge distillation to transfer the performance of large teacher models to smaller student models. This paper proposes a self-knowledge distillation method, dividing the same model into teacher and student components to bring the vision-audio capabilities of OLLMs closer to their vision-text capabilities.
@@ 8 Limitations
9 Acknowledgments
This work is supported by the National Key R&D Program of China (No. 2023YFC3310700). We sincerely appreciate the anonymous reviewers and the area chairs for their invaluable and constructive feedback, which has been instrumental in improving the quality of our study.